Match each of the following ions and molecules wit
Match each of the following ions and molecules wit
Match each of the following ions and molecules wit
Match each of the following ions and molecules wit
Answer #1

In \(\mathrm{ICl}_{2}^{-}\), the molecule has \(7+(2 \times 7)+1=22\) valence electrons. The three lone pairs on the I

atom occupy the equatorial position, the molecule acquires a linear geom etry.

In \(\mathrm{NH}_{4}^{+}\), the molecule has \(5+(1 \times 4)-1=8\) valence electrons, all of which form bonds with the \(\mathrm{H}\) atom The molecule, therefore, acquires a tetrahedral geometry.

In \(\mathrm{ClO}_{3}^{-}\), the molecule has \(7+(3 \times 6)+1=26\) valence electrons. The lon pair of electrons ar ound the \(\mathrm{Cl}\) atom repel and the molecule, therefore, acquires a trigonal pyramidal geometry.

In \(\mathrm{BCl}_{3}\), the molecule has \(3+(3 \times 7)=24\) valence electrons, all of which form bonds with the \(\mathrm{Cl}\) atom.

The molecule, therefore, acquires a trigonal planar geometry.